### v2

@@ 5 Simulation Studies
The simulation studies are divided into three part. The first part would like to numerical certify SLIDE has a superiority on empirical sample complexity with respect to $d$ and ${\gamma}$ (see Section 5.1). And the second part presented in Section 5.2 is mainly interested in another practice scenario where the number of parameters are comparable or even larger than $n$, so called high-dimensional regime. The section is closed with numerical experiments on computation performance.
@@ 5.1 Numerical Advantages on Sample Complexity
In this part, we leverages the empirical sample complexity $n_{\text{emp}}$ to evaluate and compare methods. Roughly speaking, $n_{\text{emp}}$ is empirical counterpart of $n^{*}$, recording the minimal numerical sample size that exactly recovers the underlying graphical structure. Its detailed definition is deferred to Supplementary Materials.
@@ 5.1.1 Experiments on $d$
We first investigate $n_{\text{emp}}$ with respect to $d$ to illustrate the efficacy of SLIDE. The random regular graph (RRG) topology is used in this investigation because it allows the generation of a network of any degree $d<p$ when $p,{\lambda},{\gamma}$ are all fixed. To be more specific, to disentangle the effects of $d$ from ${\lambda}$ and ${\gamma}$, we forces each node having exactly one neighbor with interaction ${\lambda}$ and the other interaction being $\beta$ such that ${\gamma}=(d-1)\beta+{\lambda}$ is a fixed constant.
Figure 1 provides a graphical representation of the corresponding Ising model in its lower right corner. A linear relationship between $n_{\text{emp}}$ and $d$ appears to exist in SLIDE (see Figure 1). This supports the $O(d)$ lower bound on $d$ obtained from our theoretical analysis of the SLIDE estimator. For the sake of comparison, we also include two benchmarked methods: (i) LASSO-regularized pseudo-likelihood estimator (RPLE) and (ii) extended LASSO (ELASSO, Van Borkulo et al., 2014). As can be seen in Figure 1, the slope of the purple line is less steep than that of the brown line, indicating that our approach is more sample-efficient than RPLE, as the rate of growth of the empirical sample size for SLIDE is lower than that for RPLE. Besides, the slope of the blue line is less steep than that of the brown line, implying ELASSO does improve PRLE in terms of sample efficiency. However, ELASSO still requires more empirical samples than SLIDE. In Supplementary Materials, we provide an additional benchmarked Ising model to show the sample complexity of RPLE/ELASSO does not grow linearly with respect to $d$ while that of SLIDE does. This implies the dominant advantage of SLIDE over RPLE/ELASSO.
Figure 1: The maximum degree $d$ ($x$-axis) versus empirical sample size ($y$-axis) scatterplot under the ferromagnetic RRGs. The network for the $d=3$ case is presented in the right-bottom corner. The ferromagnetic RRG whose couplings equal to $\beta$ and ${\lambda}$ are colored in orange and red, respectively. The three straight lines are characterized by equations $y=a+bx$, where the coefficients $a,b$ are estimated by linear regression.
@@ 5.1.2 Experiments on ${\gamma}$
Here, we empirically achieve the scale of $n_{\text{emp}}$ w.r.t. ${\gamma}$ across a variety of graphical topologies and interaction patterns via a large set of numerical tests. We take into account five benchmarked distinct Ising models considered in Lokhov et al. (2018). These Ising models cover two reference topologies including RRGs and a periodic-boundary square lattice (PBSL). The two topologies avoid the instability caused by the heterogeneity of nodes to make it simpler to extract the proper scaling with regard to ${\gamma}$. The two sorts of interactions are considered: (i) ${J}_{ij}>0$ and (ii) ${J}_{ij}\in\mathbb{R}$, where all interaction magnitudes $|{J}_{ij}|$ are set to a positive value $\beta$ except for the weakest signal ${\lambda}$. It is necessary to fix one coupling in cases (i) to ${\lambda}$ and $-{\lambda}$, respectively, and fix two couplings in cases (ii) to ${\lambda}$ and $-{\lambda}$ in order to separate the effects of ${\lambda}$ and ${\gamma}$. The combination of two topologies and two interaction types result in four Ising models. The remain one Ising model is a PBSL with interaction (i) but one of interaction is weakly negative, which is consider as a hard model for learning graphical structure (Lokhov et al., 2018).
@@
For comparison, a regularized interaction screening estimators (RISE) and its variant, logRISE, are also included in this analysis, where logRISE is a ${\gamma}$-optimal estimator (Lokhov et al., 2018). Besides, the benchmarked methods in Section 5.1.1, RPLE and ELASSO, are also taken into account. From Figure 2, the $n_{\textup{emp}}$ of five methods grows exponentially with regard to ${\gamma}$. It’s clear that the SLIDE has the finest scaling property with respect to ${\gamma}$ because its growth rate is the lowest of all these approaches across all five Ising models. Even in its worst case, the empirical scaling of SLIDE (i.e., $\exp\{3.5\}$ on PBSL with interaction (i)) is still in the optimal range with respect to the information-theoretic prediction. According to Lokhov et al. (2018), a scaling factor of 3.9 for logRISE places it in the ${\gamma}$-optimal regime, while the RISE Ising model estimator comes in at number four among the five methods under comparison. Finally, ELASSO outperforms RPLE in terms of scaling with respect to ${\gamma}$, likely due to the effectiveness of EBIC in identifying the true network structure as the sample size increases (Barber et al., 2015).
@@
Figure 2: Interactions $\beta$ ($x$-axis) versus empirical sample size ($\log_{10}$-transformed $y$-axis) scatterplot in five types of Ising models. A method with the smallest slope is the best one; otherwise, the worst. The left panels and right panels correspond to RRGs and PBSLs. The upper, middle, and bottom panels correspond to interaction types (i), (ii), and (iii), respectively. The top-left corner of each panel presents the exponent scaling $b$ of five methods, which are given by estimating the coefficients of linear models $\log(n_{\text{emp}})=a+b(d\beta)$, where $d=3$ for RRGs and $d=4$ for PBSLs.
@@ 5.2 High-dimensional Regime
TP =|$=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}\neq 0\cap{J}^{*}_{ij}\neq 0\right\}\right|,\textup{TN}=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}=0\cap{J}^{*}_{ij}=0\right\}\right|,$
FP =|$=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}\neq 0\cap{J}^{*}_{ij}=0\right\}\right|,\textup{FN}=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}=0\cap{J}^{*}_{ij}\neq 0\right\}\right|.$
@@
For each underlying graph presented in Section 5.1.2, we generate datasets with 200 observations where the dimension of binary variables varies from 18 to 34. Then, we apply our methods and the competitor on these datasets according to detailed settings shown in Supplementary Materials. Here, we also include one additional benchmarked method that leverages an upper bound on the log-partition function to simplify computation, which is referred to RLPF [^3] RLPF is not considered in Section 5.1 as it cannot exactly recover structure of Ising model when sample size is sufficiently large.. The simulation results based on 100 random replications are demonstrated in Figure 3. From the exhibited comparison, we can see that, compared to the state-of-the-art methods with $\ell_{1}$ regularization, SLIDE has a very competitive TPR, but it also would have a significantly lower FPR, which leads to its MCC, a measure that shows overall structure learning performance, would be higher than that of other approaches. Thus, from the perspective of graph support recovery, SLIDE outperforms other methods under the high-dimensional setting. The high false inclusion is not surprising because the biases caused by the $\ell_{1}$ penalty force the models to include more edges for compensation (Fan and Li, 2001; Xue et al., 2012).
@@ 5.3 Runtime Analysis
Adopting the numerical settings in Section 5.2, we further investigate the computational performance in these cases. As shown in Supplementary Materials, among all nodewise methods, the SLIDE demonstrates the fastest computational speed on five benchmarked Ising models, followed by ELASSO, logRISE, RISE, and finally RPLE method. Remarkably, the computational advantage is tremendous, with up to a 100-fold speedup compared to RPLE/RISE/logRISE and 10-fold speedup compared to ELASSO. Our empirical results refute the widely held belief that the best-subset selection approach is not feasible for learning high-dimensional Ising models. More importantly, the fitted curves (displayed in the top-left corner of each panel) depicting dimensionality against computation time show that the runtime of SLIDE scales quadratically with respect to $p$. This confirms the computational complexity of SLIDE is polynomial. Finally, RLPF is as competitive as SLIDE when $p$ is small, and it marginally outperforms SLIDE when $p$ is large. However, this slight computational advantage of RLPF is not adequately attractive, as RLPF cannot exactly reconstruct some Ising models.
@@ 6 US Senate Voting Data Analysis
From Figure 4, the 112th and 117th Congresses were visually more distinct. The 112th Congress had a greater degree of partisanship than any other. The Washington Post [^5] https://www.washingtonpost.com/news/wonk/wp/2013/01/17/its-official-the-112th-congress-was-the-most-polarized-ever/ reported that the 112th Congress was the most polarized ever as measured by the DW-NOMINATE score (Poole and Rosenthal, 1985). We apply the spectral clustering method (Shi and Malik, 2000) on the reconstructed Ising model to make predictions, and get an unsupervised classifier with an accuracy rate of nearly $90\%$. Next, the Democratic senators in the 117th Congress displayed the most concentrated and out-of-the-ordinary pattern. This Senate is made up of 50 Republicans, 48 Democrats, and 2 Independents who caucus with the Democrats. This makes the Senate effectively evenly split, only the third time in US history. Without any leeway for dissent, to support the agendas of the Democrat president, the Democrats need to be fully united [^6] https://fivethirtyeight.com/features/why-house-democrats-may-be-more-united-than-they-seem/.
@@ 7 Conclusion and Discussion
More efforts are deserves to enlarge the impact of sparse learning approach and the algorithm for solving. For example, one promising direction is extending our approach to recover Potts models for discrete variables (Wu, 1982) and Ising model with random effects (Fan and Li, 2012; He et al., 2023a, b). From application, our solution facilitates the new findings on cutting-edge scientific problems such as functional connectome and disease connectome.

### v3

@@ 5 Simulation Studies
The simulation studies are divided into three part. The first part would like to certify SLIDE has a superiority on empirical sample complexity $n^{\textup{emp}}$ with respect to $d$ and $\gamma$} (see Section 5.1). And the second part presented in Section 5.2 is mainly interested in another practice scenario where the number of parameters are comparable or even larger than $n$, so called high-dimensional regime. The section is closed with numerical experiments on computation performance.
@@ 5.1 Numerical Advantages on Sample Complexity
In this part, we leverage the empirical sample complexity $n^{\textup{emp}}$ to evaluate and compare methods. Roughly speaking, $n^{\textup{emp}}$ is empirical counterpart of $n^{*}$, recording the minimal numerical sample size that exactly recovers the underlying graphical structure. Its detailed definition is deferred to the Supplementary Materials. We also empirically verify $n^{\textup{emp}}$ is proportional to $\log p$ and ${\lambda}^{-2}$, which is matched to the sample complexity of SLIDE. Please see the Supplementary Materials for detailed results.
@@ 5.1.1 Experiments on $d$
We first investigate $n^{\textup{emp}}$ with respect to $d$ to illustrate the efficacy of SLIDE. The random regular graph (RRG) topology is used in this investigation because it allows the generation of a network of any degree $d<p$ when $p,{\lambda},{\gamma}$ are all fixed. To be more specific, to disentangle the effects of $d$ from $\lambda$} and $\gamma$}, we force each node having exactly one neighbor with interaction $\lambda$} and the other interaction being $\beta$ such that ${\gamma}=(d-1)\beta+{\lambda}$ is a fixed constant.
Figure 1 provides a graphical representation of the corresponding Ising model in its lower right corner. A linear relationship between $n^{\textup{emp}}$ and $d$ appears to exist in SLIDE (see Figure 1), with adjusted $R^{2}$ reaches 0.97. This supports the $O(d)$ lower bound on $d$ obtained from our theoretical analysis of the SLIDE estimator. For the sake of comparison, we also include two benchmarked methods: (i) LASSO-regularized pseudo-likelihood estimator (RPLE) and (ii) extended LASSO (ELASSO, Van Borkulo et al., 2014). As can be seen in Figure 1, the slope of the purple line is less steep than that of the brown line, indicating that our approach is more sample-efficient than RPLE, as the rate of growth of the empirical sample size for SLIDE is lower than that for RPLE. Besides, the slope of the blue line is less steep than that of the brown line, implying ELASSO does improve PRLE in terms of sample efficiency. However, ELASSO still requires more empirical samples than SLIDE. In the Supplementary Materials, we provide an additional benchmarked Ising model to show the sample complexity of RPLE/ELASSO does not grow linearly with respect to $d$ while that of SLIDE does. This implies the dominant advantage of SLIDE over RPLE/ELASSO.
Figure 1: The maximum degree $d$ ($x$-axis) versus empirical sample size ($y$-axis) scatterplot under the ferromagnetic RRGs. The three straight lines are characterized by equations $y=a+bx$, where the coefficients $a,b$ are estimated by linear regression. The adjusted $R^{2}$ of each fitted linear model is reported in the left-top corner.
@@ 5.1.2 Experiments on $\gamma$}
Here, we empirically achieve the scale of $n^{\textup{emp}}$ w.r.t. $\gamma$} across a variety of graphical topologies and interaction patterns via a large set of numerical tests. We take into account five benchmarked distinct Ising models considered in Lokhov et al. (2018). These Ising models cover two reference topologies including RRGs and a periodic-boundary square lattice (PBSL). The two topologies avoid the instability caused by the heterogeneity of nodes to make it simpler to extract the proper scaling with regard to $\gamma$}. The two sorts of interactions are considered: (i) ${J}_{ij}>0$ and (ii) ${J}_{ij}\in\mathbb{R}$, where all interaction magnitudes $|{J}_{ij}|$ are set to a positive value $\beta$ except for the weakest signal $\lambda$}. It is necessary to fix one coupling in cases (i) to $\lambda$} and $\lambda$}, respectively, and fix two couplings in cases (ii) to $\lambda$} and $\lambda$} in order to separate the effects of $\lambda$} and $\gamma$}. The combination of two topologies and two interaction types result in four Ising models. The remain one Ising model is a PBSL with interaction (i) but one of interaction is weakly negative, which is considered as a hard model for learning graphical structure (Lokhov et al., 2018).
@@
For comparison, a regularized interaction screening estimators (RISE) and its variant, logRISE, are also included in this analysis, where logRISE is a $\gamma$}-optimal estimator (Lokhov et al., 2018). Besides, the benchmarked methods in Section 5.1.1, RPLE and ELASSO, are also taken into account. From Figure 2, the $n_{\textup{emp}}$ of five methods grows exponentially with regard to $\gamma$} (as evidence by large adjusted $R^{2}$ values). It’s clear that the SLIDE has the finest scaling property with respect to $\gamma$} because its growth rate is the lowest of all these approaches across all five Ising models. Even in its worst case, the empirical scaling of SLIDE (i.e., $\exp\{3.5\}$ on PBSL with interaction (i)) is still in the optimal range with respect to the information-theoretic prediction. According to Lokhov et al. (2018), a scaling factor of 3.9 for logRISE places it in the $\gamma$}-optimal regime, while the RISE Ising model estimator comes in at number four among the five methods under comparison. Finally, ELASSO outperforms RPLE in terms of scaling with respect to $\gamma$}, likely due to the effectiveness of EBIC in identifying the true network structure as the sample size increases (Barber et al., 2015).
@@
Figure 2: Interactions $\beta$ ($x$-axis) versus empirical sample size ($\log_{10}$-transformed $y$-axis) scatterplot in five types of Ising models. A method with the smallest slope is the best one; otherwise, the worst. The left panels and right panels correspond to RRGs and PBSLs. The scatterplots in upper, middle, and bottom panels correspond to interaction types (i), (ii), and (iii), respectively. The top-left corner of each panel presents the exponent scaling $b$ of five methods, which are given by estimating the coefficients of linear models $\log(n^{\textup{emp}})=a+b(d\beta)$, where $d=3$ for RRGs and $d=4$ for PBSLs. The table in the right-bottom corner reports the adjusted $R^{2}$ of the fitted linear models.
@@ 5.2 High-dimensional Regime
TP $=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}\neq 0\cap{J}^{*}_{ij}\neq 0\right\}\right|,\textup{TN}=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}=0\cap{J}^{*}_{ij}=0\right\}\right|,$
FP $=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}\neq 0\cap{J}^{*}_{ij}=0\right\}\right|,\textup{FN}=\left|\left\{(i,j)\mid\widehat{{J}}_{ij}=0\cap{J}^{*}_{ij}\neq 0\right\}\right|.$
@@
For each underlying graph presented in Section 5.1.2, we generate datasets with 200 observations where the dimension of binary variables varies from 18 to 34. Then, we apply our methods and the competitor on these datasets according to detailed settings shown in the Supplementary Materials. Here, we also include one additional benchmarked method that leverages an upper bound on the log-partition function to simplify computation, which is referred to RLPF [^4] RLPF is not considered in Section 5.1 as it cannot exactly recover structure of Ising model when sample size is sufficiently large (see the Supplementary Materials).. The simulation results based on 100 random replications are demonstrated in Figure 3. From the exhibited comparison, we can see that, compared to the state-of-the-art methods with $\ell_{1}$ regularization, SLIDE has a very competitive TPR, but it also would have a significantly lower FPR, which leads to its MCC, a measure that shows overall structure learning performance, would be higher than that of other approaches. Thus, from the perspective of graph support recovery, SLIDE outperforms other methods under the high-dimensional setting. The high false inclusion is not surprising because the biases caused by the $\ell_{1}$ penalty force the models to include more edges for compensation (Fan & Li, 2001; Xue et al., 2012).
@@ 5.3 Runtime Analysis
Adopting the numerical settings in Section 5.2, we further investigate the computational performance in these cases. As shown in the Supplementary Materials, among all nodewise methods, the SLIDE demonstrates the fastest computational speed on five benchmarked Ising models, followed by ELASSO, logRISE, RISE, and finally RPLE method. Remarkably, the computational advantage is tremendous, with up to a 100-fold speedup compared to RPLE/RISE/logRISE and 10-fold speedup compared to ELASSO. The results in the Supplementary Materials refute the widely held belief that the best-subset selection approach is not feasible for learning high-dimensional Ising models. More importantly, the fitted curves (displayed in the top-left corner of each panel) depicting dimensionality against computation time show that the runtime of SLIDE scales quadratically with respect to $p$. This confirms the computational complexity of SLIDE is polynomial. Finally, RLPF is as competitive as SLIDE when $p$ is small, and it marginally outperforms SLIDE when $p$ is large. However, this slight computational advantage of RLPF is not adequately attractive, as RLPF cannot exactly reconstruct some Ising models.
@@ 6 US Senate Voting Data Analysis
From Figure 4, the 112th and 117th Congresses were visually more distinct. The 112th Congress had a greater degree of partisanship than any other. The Washington Post [^6] https://www.washingtonpost.com/news/wonk/wp/2013/01/17/its-official-the-112th-congress-was-the-most-polarized-ever/ reported that the 112th Congress was the most polarized ever as measured by the DW-NOMINATE score (Poole & Rosenthal, 1985). We apply the spectral clustering method (Shi & Malik, 2000) on the reconstructed Ising model to make predictions, and get an unsupervised classifier with an accuracy rate of nearly $90\%$. Next, the Democratic senators in the 117th Congress displayed the most concentrated and out-of-the-ordinary pattern. This Senate is made up of 50 Republicans, 48 Democrats, and 2 Independents who caucus with the Democrats. This makes the Senate effectively evenly split, only the third time in US history. Without any leeway for dissent, to support the agendas of the Democrat president, the Democrats need to be fully united [^7] https://fivethirtyeight.com/features/why-house-democrats-may-be-more-united-than-they-seem/.
@@ 7 Conclusion and Discussion
More efforts are deserves to enlarge the impact of sparse learning approach and the algorithm for solving. For example, one promising direction is extending our approach to recover Potts models for discrete variables (Wu, 1982) and Ising model with random effects (Fan & Li, 2012; He et al., 2023a, b). From application, our solution facilitates the new findings on cutting-edge scientific problems such as functional connectome and disease connectome.